OpenStudy (anonymous):

I am having a hard time understanding the way much of the online math resources are written. They all seem to be using a very mathematical sort of dialect (thus we can see... or to find x we need to complete a square, with no word on how one might accomplish this "completion of a square"), which in my opinion is counterproductive--the very people needing help with math are the very ones who won't understand the math language. Anyways, that's a bit long-winded, so here's the problem: "If a parabola opening to the left has a vertex at (-1, -3) and passes through point (-3,-2), what is the co

OpenStudy (anonymous):

Um, you chopped off a bit of the question, there....

OpenStudy (anonymous):

Right, thanks. Not sure what happened, is there a space limit here? It finishes like this "coefficient of the squared term in the parabola's equation?"

OpenStudy (anonymous):

Suppose we are talking about a parabola of form y^2 = 4ax (vertex at (0,0) then (y-a) ^2 = 4k(x-b)^2 where vertex is at (a,b) so (y+3)^2 = 4 k(x+1) then find k using the given point and rearrange in the form x = f(y)

OpenStudy (anonymous):

What you've put there is quite different than the way my math book has it; my book does not have a k or an f for porabolas. What are they, exactly? My book has a and b, so I'll assume they're the same?

OpenStudy (anonymous):

f(y) just means "a function of y" , k is a constant (use c or whatever you like best) (I have used a and b as stand ins for the vertex)